Please help me!!!
how many pounds of hamburger that costs $1.10 per pound must be mixed with 50 pounds of hamburger that costs $1.80 per pound to make a mixture that costs $1.30 per pound?

Answer by Paul(988)

1.10x+1.80(50-x)=1.30(50)
1.10+90-1.80x=65
-0.7x=-25
x=35(rounded)